a) Solve (𝐷


2 + 1)𝑦 = 3π‘₯ βˆ’ 8 cot π‘₯ where 𝐷 =


𝑑


𝑑π‘₯



b) Find orthogonal trajectory to the curve given by π‘Ÿ = π‘Ž(1 + cos πœƒ)

1
Expert's answer
2022-01-04T07:29:40-0500

Solution:

a):

"(D^2+1)y=3x-8\cot x"

Its auxiliary equation is "m^2+1=0"

"\Rightarrow m=\pm1"

The complementary function is: "y_c(x)=c_1\cos x+c_2\sin x"

Let "u(x)=\cos x,v(x)=\sin x"

Now, particular integral: "y_p(x)=A.u(x)+B.v(x)"

We use the method of parameters.

"A=-\int \dfrac{u(x)Q(x)}{w(u,v)}dx,B=\int \dfrac{v(x)Q(X)}{w(u,v)}dx
\\w(u,v)=\begin{vmatrix} u&v\\u'&v'\end{vmatrix}=\begin{vmatrix} \cos x&\sin x\\-\sin x&\cos x\end{vmatrix}=1"

And "Q(x)=3x-8\cot x"

Then, "A=-\int \dfrac{\cos x(3x-8\cot x)}{1}dx"

"=-3\int x\cos xdx+8\int\cos x\cot xdx
\\=-3[x \sin x-1\times -\cos x]+8[\ln \tan(x/2)+\cos x]
\\=-3x\sin x+5\cos x+8 \ln \tan(x/2)" [Using integration by parts]

Next, "B=\int \dfrac{\sin x(3x-8\cot x)}{1}dx"

"=3\int x\sin xdx-8\int\sin x\cot xdx
\\=3[x\cos x-1\times-\sin x]-8\sin x
\\=-3x\cos x-5\sin x" [Using integration by parts]

So, the particular integral is

"y_p(x)=A.u(x)+B.v(x)
\\=(-3x\sin x+5\cos x+8 \ln \tan(x/2))(\cos x)+(-3x\cos x-5\sin x)(\sin x)
\\=-6x\sin x \cos x+5[\cos^2x-\sin^2x]+8\ln \tan(x/2)
\\=-6x\sin x \cos x+5\cos2x+8\ln \tan(x/2)"

Thus, the solution is

"y=y_c(x)+y_p(x)
\\y=c_1\cos x+c_2\sin x-6x\sin x \cos x+5\cos2x+8\ln \tan(x/2)"

b):

Given curve is "r=a(1+\cos \theta)" ...(i)

Differentiating it, "d r+a \sin \theta d \theta=0" ...(ii)

On eliminating the parameter ' a ' we get,

"d r+\frac{r \sin \theta}{1+\cos \theta} d \theta=0 \quad or \quad \frac{-1}{r} \frac{d r}{d \theta}=\frac{\sin \theta}{1+\cos \theta}" ...(iii)

This is the differential equation of the direction field "D_{1}" for the given family "F_{1}" .

To find the differential equation of the direction field orthological to "D_{1}", we replace "\frac{d r}{d \theta}"

by "-r^{2} \frac{d \theta}{d r}" as evident from the geometry.

Where "\phi_{2}\ and\ \Psi_{2}" are the angles for the orthogonal trajectory corresponding to the angles "\phi_{1}\ and\ \Psi_{1}" for the given family (Fig. 1) as shown below.


"\begin{aligned}

\text { Clearly, } \quad \phi_{2} &=\left(\phi_{1} \pm \frac{\pi}{2}\right) \text {, whence } \tan \phi_{2}=\tan \left(\phi_{1} \pm \frac{\pi}{2}\right)=-\cot \phi_{1} \\

\left(r \frac{d \theta}{d r}\right)_{2} &=-\left(\frac{1}{r} \frac{d r}{d \theta}\right)_{1}

\end{aligned}"

Whence from (3), the differential equation for the orthogonal trajectories becomes

"r \frac{d \theta}{d r}=\frac{\sin \theta}{1+\cos \theta}=\tan \frac{\theta}{2}"

which is a case of variable-separable, and on integration gives

"\log r=2 \log \sin \frac{\theta}{2}+\log 2 c"

This is the equation of the orthogonal family "F_{2}" .

Since "r=c(1-\cos \theta)" represents the same curve as "r=c(1+\cos \theta)," the member of

"F_{2}" with label c is the same as the member of "F_{1}" with label "a=-c" . Thus the given family is selforthogonal.